OpenStudy (anonymous):

Solve the equation 3.75x + 3.7 = 1.7 + 1.75x

OpenStudy (anonymous):

Combine like terms.

OpenStudy (anonymous):

how

OpenStudy (shaik0124):

subtract -1.75x on both sises 3.75x-1.75x+3.7=1.7+1.75x-1.75x 2.0x+3.7=1.7 subtrct -3.7 on both sides 2.0x+3.7-3.7=1.7-3.7 2.0x=-2.0 divide 2.0 on both sides x=-1